# Semi-parametric bulk and tail regression using spline-based neural networks

### v1

@@ 4.2 Results
Figure 6 provides estimates of the variable importance scores for both the SPQR and xSPQR model fits. We consider three time periods: the two test periods, 1990–1994 and 2016–200, and the training period, 1995-2015. Across both models and all time periods, minimum relative humidity rmin is identified as the most important variable for predicting the distribution of U.S. wildfire burnt areas across all quantile levels, $\tau$. Additionally, while precipitation appears less important than maximum temperature (tmax) in the period 1990–1994, both current (pr_curr) and previous year (pr_prev) precipitation show significant increases in relative importance of these variables between 1995 and 2020, with both overtaking maximum temperature during this period. The covariates windspeed (wspd) and year (fire_yr) remain relatively unimportant throughout the entire time period. This suggests that dryness is a bigger driving factor of burnt area than high temperatures: this is not surprising, as while high temperatures increase ignition rates (from natural causes) and windspeed affects the rate of spread, dryness is directly tied to the total area that has potential for burning. Figure 6 also illustrates an overall increase in the VI scores over time, suggesting that these meteorological variables may be becoming increasingly more important. Table 2 provides the variable importance scores for the shape parameter $\xi$, for each of the three time periods. The most important variable for estimating $\xi$ is the previous years’ precipitation, pr_prev. Our interpretation is that while low humidity leads to more burnt area (as rmin is in the top three most important variables for the tail), the most devastating wildfires, in terms of burnt area, depend more on whether the region has been dry for prolonged periods of time. On the other hand, the year is estimated to be largely inconsequential for determining the tail of the conditional density. Beyond this, the results are largely consistent with the relative variable importance analysis for the body of the distribution. Overall, across the body and the tails, we identify dryness as having the largest effect on the distribution of U.S. wildfire burnt areas.
@@
Table 2: Estimates of the variable importance scores for the shape parameter, $\xi(\mathbf{x})$, of the xSPQR model fit.
@@
Time period pr_prev rmin tmax wspd pr_curr fire_yr
1990–1994 2.35 2.08 1.41 0.90 0.91 0.22
1995–2015 2.87 2.13 1.82 1.72 1.21 0.61
2016–2020 2.39 1.69 1.42 2.26 1.58 0.50
@@
Figure 7 provides the density of $\xi(\mathbf{x})$ estimates for the training period, 1995–2015, and the two testing periods of 1990–1994 and 2016–2020. Figure 7 illustrates a positive temporal trend in the estimates. The change is most noticeable between 1990–1994 and 1995–2015, and may suggest that U.S. wildfires are becoming more heavy-tailed, i.e., extreme, over time. While the estimates for 2016–2020 are more concentrated compared to the training period, they are both centered around approximately the same value. We note that the response data have undergone a square-root transformation, and a value of $\xi(\mathbf{x})>0.25$ on the response scale would translate to approximately $\xi(\mathbf{x})>0.5$ on the original scale; in the latter case, the blended GP density does not have finite variance, which suggests that the original response data are extremely heavy-tailed.
@@
Figure 8 illustrates the spatial variability in the estimated quantiles of the burnt area response. For each fire in the dataset, we estimate the median and $0.999$-quantile using the fitted xSPQR model; the former quantity is estimated via the SPQR model for the bulk while the latter is estimated via the GP model for the upper tail. The quantile is squared to obtain the actual burnt area in 1000s of acres, and then aggregated over Level III (L3) ecoregions. The ecoregions group ecosystems that are similar in terms of their biotic, abiotic, terrestrial, and aquatic ecosystem components, including human activity (McMahon et al., , 2001; Omernik and Griffith, , 2014); ecoregion boundaries are made available by the U.S. Environmental Protection Agency [^1] https://www.epa.gov/eco-research/level-iii-and-iv-ecoregions-continental-united-states, with lower levels indicating more aggregation (and therefore larger ecoregions). From Figure 8, estimates of the quantiles show spatial variability across the country, with higher values concentrated in the western regions of the U.S. The illustrated east-west gradient is more pronounced in the map of the estimated 0.999-quantile than the median; alongside the differences in the relative VI scores for the bulk and tail (see Figure 6 and Table 2), the differences in the spatial patterns of the median and tail quantiles provides further evidence that the covariates have different impacts on the bulk and upper-tail of the wildfire distribution.
@@
Figure 9: Estimated conditional log-density functions $\log f(y|\mbox{\bf x})$ for the two largest wildfires in the dataset. The response variable $Y$ is the square-root burnt area, which were originally measured in 1000s of acres. The red lines denote the observed value for the fires, while the dashed black lines represent the largest value that is observed in the training data.
@@
Finally, we focus on the two most extreme fires in the dataset (in terms of burnt area). We refer to them as the Klamath mountains and Southwestern tablelands fires, based on their L3 ecoregion of occurrence (Lawler and Shaby, , 2024). As mentioned above, they are both present in the second test period, 2016–2020, and are therefore outside the range of the training data used to fit our models. Both fires were associated with moderate temperatures ($<30\degree$C), but with very low humidity ($<25\%$). Additionally, in the case of the Klamath mountains fire, the aggregate precipitation over the past month (pr_curr) was in the lowest $0.07$ quantile of the training data. Figure 9 provides the xSPQR estimates of the conditional log-density for these two fires. In each case, the observed wildfires are well beyond the observed maximum of the training data, and are modelled explicitly by the GP upper-tail of the xSPQR model. The estimated shape parameters for these two fires, while high, are lower than the median and $0.8$ quantile, respectively, of the entire test set of $\xi$ predictions. Figure 9 highlights the flexibility of the xSPQR model; unlike existing fully-parametric bulk and tail models (see Section 1), xSPQR is able to capture both the heavy-tails and high kurtosis exhibited by the data.
@@ 6 Discussion
In this work, we developed a flexible framework for conditional bulk and tail estimation called extremal semi-parametric quantile regression (xSPQR). To achieve this, we propose the blended generalised Pareto (bGP) distribution, which is a new univariate model for the entire range of data that is EVT-compliant, in that its upper-tail follows exactly a GP distribution. We use the bGP distribution to combine a deep parametric GP regression model with the traditional SPQR framework for conditional density estimation. The flexible nature of SPQR ensures that parametric assumptions do not need to be made about the bulk of the conditional distribution, which can instead be arbitrarily flexible. The xSPQR framework addresses one of the main weaknesses of SPQR by allowing the modelling of heavy-tailed distributions and making predictions outside the range of the observed data. We also obtain variable importance estimates for xSPQR using accumulated local effects; as xSPQR allows covariates to separately affect the bulk and the tail of the distribution, we can perform inference on variable importance for the tails separately of that for the bulk. A simulation study and a real-data application to U.S. wildfire burnt areas demonstrated the efficacy of xSPQR.
@@
In our current work, the hyper-parameters of the blending function, $(p_{a},p_{b},c_{1},c_{2}),$ are carefully chosen using numerical experiments and via grid search algorithms. Arbitrarily choosing these hyper-parameters can lead to pathological cases where the density function, $h(y|\mathcal{W},\xi)$ in (6), is negative for $y$ in the blending interval, $[a,b]$. In practice, we found that this phenomenon was extremely rare, and the risk of it occurring during training of the xSPQR model was mitigated by the use of model pre-training (Section 3.4); however, future work will investigate regularization techniques for fully precluding this phenomenon.
In this work, we only considered estimation of heavy upper-tails, i.e., $\xi(\mathbf{x})>0$; this constraint was imposed to prevent numerical problems when fitting the xSPQR model. While theoretically xSPQR can handle $\xi(\mathbf{x})<0$, and we thus presented the blended GP in Section 3.2 for general $\xi\in\mathbb{R}$, training of xSPQR in this case requires a much more involved numerical optimisation scheme or reparameterisation of the model; see, e.g., Pasche and Engelke, (2024) and Mackay et al., (2024). We leave such a consideration as future work.
@@
Another simple extension of the xSPQR framework is to allow for different subsets of variables to affect the bulk and upper-tail of the distribution, which can be achieved by modelling $\mathcal{W}(\mathbf{x})$ and $\xi(\mathbf{x})$ in (7) using separate neural networks (with separate inputs). Further work may also consider a blended GP distribution and resulting xSPQR model that allows for extrapolation in both the lower- and upper-tail of the conditional distribution. An extension of (4), which blends a constituent model for the bulk with two independent GP distributions for the lower- and upper-tail, is trivial; however, inference for such a model using deep learning may be numerically difficult.
Declarations
Ethical Approval
Not Applicable
Availability of supporting data
Competing interests
The authors have no relevant financial or non-financial interests to disclose.
@@ Appendix A Construction of $M$- and $I$-splines
[…]
@@
In the case of SPQR, we use an order of $d=3$ for the polynomial basis. We note that $B$-splines are also closely related to $M$-splines, as the $B$-spline basis function is defined by $B_{k}=(t_{k+d}-t_{k})M_{k}/d$ (Ramsay, , 1988).
Ramsay, (1988) also introduced integrated splines, or $I$-splines, over $[L,U]$. These are built using basis functions

### v2

@@ 4.2 Results
Figure 6 provides estimates of the variable importance scores for both the SPQR and SPQRx model fits. We consider three time periods: the two test periods, 1990–1994 and 2016–200, and the training period, 1995-2015. Across both models and all time periods, minimum relative humidity rmin is identified as the most important variable for predicting the distribution of U.S. wildfire burnt areas across all quantile levels, $\tau$. Additionally, while precipitation appears less important than maximum temperature (tmax) in the period 1990–1994, both current (pr_curr) and previous year (pr_prev) precipitation show significant increases in relative importance of these variables between 1995 and 2020, with both overtaking maximum temperature during this period. The covariates windspeed (wspd) and year (fire_yr) remain relatively unimportant throughout the entire time period. This suggests that dryness is a bigger driving factor of burnt area than high temperatures: this is not surprising, as while high temperatures increase ignition rates (from natural causes) and windspeed affects spread, dryness is more directly tied to the total area that has potential for burning. Figure 6 also illustrates an overall increase in the VI scores over time, suggesting that these meteorological variables may be becoming increasingly more important. Table 2 provides the variable importance scores for the shape parameter $\xi$, for each of the three time periods. The importance scores for the tail largely follow the same trends as the importance scores for the body of the distribution, with short-term and long-term dryness dominating, and tmax showing slight decreases. In particular, pr_prev shows an eight-fold increase between the 1990–1994 and 2016–2020 periods. Overall, across the body and the tails, we identify dryness as having the largest effect on the distribution of U.S. wildfire burnt areas.
@@
Table 2: Estimates of the variable importance scores for the shape parameter, $\xi(\mathbf{x})$, of the SPQRx model fit.
@@
Time period rmin pr_curr pr_prev tmax wspd fire_yr
1990–1994 2.13 0.83 0.38 1.79 0.77 0.17
1995–2015 2.74 1.53 1.95 1.72 0.23 0.55
2016–2020 3.29 2.35 3.01 1.17 0.45 0.04
@@
Figure 7 provides the density of $\xi(\mathbf{x})$ estimates for the training period, 1995–2015, and the two testing periods of 1990–1994 and 2016–2020. Figure 7 illustrates a positive temporal trend in the estimates, and suggests that U.S. wildfires are becoming more heavy-tailed, i.e., extreme, over time. We note that the response data have undergone a square-root transformation, and a value of $\xi(\mathbf{x})>0.25$ on the response scale would translate to approximately $\xi(\mathbf{x})>0.5$ on the original scale. In the latter case, the blended GP density does not have finite variance, which suggests that the original response data are extremely heavy-tailed.
@@
Figure 8 illustrates the spatial variability in the estimated quantiles of the burnt area response. For each fire in the dataset, we estimate the median and $0.999$-quantile using the fitted SPQRx model. The former quantity is estimated via the SPQR model for the bulk while the latter is estimated via the GP model for the upper-tail. The quantile is squared to obtain the actual burnt area in 1000s of acres, and then aggregated over Level III (L3) ecoregions. The ecoregions group ecosystems that are similar in terms of their biotic, abiotic, terrestrial, and aquatic ecosystem components, including human activity (McMahon et al., 2001; Omernik and Griffith, 2014). Ecoregion boundaries are made available by the U.S. Environmental Protection Agency [^2] https://www.epa.gov/eco-research/level-iii-and-iv-ecoregions-continental-united-states, with lower levels indicating more aggregation (and therefore larger ecoregions). From Figure 8, estimates of the quantiles show spatial variability across the country, with higher values concentrated in the western regions of the U.S. The illustrated east-west gradient is more pronounced in the map of the estimated 0.999-quantile than the median. Alongside the differences in the relative VI scores for the bulk and tail (see Figure 6 and Table 2), the differences in the spatial patterns of the median and tail quantiles provides further evidence that the covariates have different impacts on the bulk and upper-tail of the wildfire distribution.
@@
Figure 9: Estimated conditional log-density functions $\log f(y|\mbox{\bf x})$ for the two largest wildfires in the dataset. The response variable $Y$ is the square-root of burnt area, originally measured in 1000s of acres. The red lines denote the observed value for the fires, while the dashed blue lines represent the largest value that is observed in the training data.
Figure 10: Bootstrapped estimates of the exceedance probability $\Pr(Y>y|\mbox{\bf x})$ for the two largest wildfires in the dataset. The response variable $Y$ is the square-root of burnt area, originally measured in 1000s of acres. The red lines denote the tail probability estimated by SPQRx, while the dashed black lines represent $95\%$ confidence intervals based on 200 bootstrap samples.
Figure 11: Bootstrapped estimates of the shape parameter $\xi$ for the two largest wildfires in the dataset. The red lines denote the estimate from SPQRx, while the dashed black lines represent $95\%$ confidence intervals based on 200 bootstrap samples.
@@
Finally, we focus on the two most extreme fires in the dataset (in terms of burnt area). We refer to them as the Klamath mountains and Southwestern tablelands fires, based on their L3 ecoregion of occurrence (Lawler and Shaby, 2024). As mentioned above, they are both present in the second test period, 2016–2020, and are therefore outside the range of the training data used to fit our models. Both fires were associated with moderate temperatures ($30\degreeC$), but with very low relative humidity ($25\%$). Additionally, in the case of the Klamath mountains fire, the aggregate precipitation over the past month (pr_curr) was in the lowest $0.07$ quantile of the training data. Figure 9 provides the SPQRx estimates of the conditional log-density for these two fires. In each case, the observed wildfires are well beyond the observed maximum of the training data, and are modelled explicitly by the GP upper-tail of the SPQRx model. Figure 9 highlights the flexibility of the SPQRx model which, unlike existing fully-parametric bulk and tail models (see Section 1), is able to capture both the heavy-tails and high kurtosis exhibited by the data. Figure 10 plots a histogram of the estimated exceedance probability $\Pr(Y>y|\mbox{\bf x})$ for the two fires. The Klamath mountain fire has an estimated exceedance probability of $5.56\times 10^{-5}$ and a $95\%$ confidence interval of $(9.15\times 10^{-6},1.74\times 10^{-4})$. The Southwestern tablelands fire has an exceedance probability of $0.0003$ with a $95\%$ confidence interval of $(0.0001,0.0004)$. Finally, Figure 11 plots uncertainty estimates of $\xi$ for the two fires: the Klamath mountains fire ($\hat{\xi}=0.12$) has a $95\%$ interval of $(0.02,0.32)$, while the Southwestern tablelands fire ($\hat{\xi}=0.19$) has a 95% confidence interval of ($0.06,0.34$). We note that the larger estimate of $\xi(\mbox{\bf x})$ among these two is actually associated with the smaller wildfire. This demonstrates the necessity of a modelling approach like SPQRx that is adept at capturing how covariates affect both the bulk and tail of the burnt area distribution.
@@ 6 Discussion
In this work, we developed a flexible framework called semi-parametric quantile regression for extremes (SPQRx) for conditional bulk and tail estimation. To achieve this, we propose the blended generalised Pareto (bGP) distribution, which is a new univariate model for the entire range of data that is EVT-compliant, in that its upper-tail follows exactly a GP distribution. We use the bGP distribution to combine a deep parametric GP regression model with the traditional SPQR framework for conditional density estimation. The flexible nature of SPQR ensures that parametric assumptions do not need to be made about the bulk of the conditional distribution, which can instead be arbitrarily flexible. The SPQRx framework addresses one of the main weaknesses of SPQR by allowing the modelling of heavy-tailed distributions and making predictions outside the range of the observed data. We also obtain variable importance estimates for SPQRx using accumulated local effects. As SPQRx allows covariates to separately affect the bulk and the tail of the distribution, we can perform inference on variable importance for the tails separately of that for the bulk. A simulation study and a real-data application to U.S. wildfire burnt areas demonstrated the efficacy of SPQRx.
@@
In our current work, the hyper-parameters of the blending function, $(p_{a},p_{b},c_{1},c_{2}),$ are carefully chosen using numerical experiments and via grid search algorithms. Arbitrarily choosing these hyper-parameters can lead to pathological cases where the density function, $h(y|\mathcal{W},\xi)$ in (6), is negative for $y$ in the blending interval, $[a,b]$ (see Section 3.2). In practice, we found that this phenomenon was extremely rare, and we mitigated the risk of it occurring during training of the SPQRx model by pre-training (Section 3.4) and penalisation (Appendix B). Future work may investigate the specific constraints required on the constituent distribution functions and weighting function that guarantee validity of the resulting blended GP distribution.
@@
Another simple extension of the SPQRx framework is to allow for different subsets of variables to affect the bulk and upper-tail of the distribution, which can be achieved by modelling $\mathcal{W}(\mathbf{x})$ and $\xi(\mathbf{x})$ in (7) using separate neural networks (with separate inputs). Further work may also consider a blended GP distribution and resulting SPQRx model that allows for extrapolation in both the lower- and upper-tail of the conditional distribution. An extension of (4), which blends a constituent model for the bulk with two independent GP distributions for the lower- and upper-tail, is trivial. However, inference for such a model using deep learning may be numerically difficult.
As noted in Section 2.2, the SPQR representation in (1) can approximate any continuous density function as the number $K$ of basis functions goes to infinity. Further work could investigate similar theoretical results for SPQRx. As we have equality between the SPQRx and SPQR density and distributions functions on the compact support $[0,a],$ for $a$ the $p_{a}$-quantile of the SPQR distribution, we might expect a weaker approximation result to hold for SPQRx. A particular case of interest might be that where $p_{a}\rightarrow 1$, and the SPQRx distribution function converges to its SPQR constituent distribution.
Finally, extending SPQRx to a multivariate setting would allow us to study non-stationarity in multivariate distributions, whilst also providing a flexible model which is appropriate for the joint bulk and tails. This would require describing a copula for the bulk using, e.g., tensor splines (Zong and Lam, 1998), alongside an extremal bivariate copula or blended GP marginal tails. Such an approach may provide an alternative to the parametric multivariate bulk and tail models of, e.g., André et al. (2024) and Alotaibi et al. (2025). An alternative approach could be to couch SPQRx within the angular-radial modelling frameworks of, e.g., Mackay et al. (2025) or Murphy-Barltrop et al. (2025).
@@ Appendix A Construction of $M$- and $I$-splines
A polynomial spline is a piecewise polynomial function over a compact interval, say $[L,U],$ with the constituent polynomials having the same degree and connecting smoothly at predetermined knots. Several classes of splines, e.g., $B$-splines, also constitute basis functions of their function space (Prautzsch et al., 2002). Similarly, the $M$-spline family of basis functions, $\{M_{k}\}_{k=1}^{K},$ is suited to representing densities (Curry and Schoenberg, 1966). The $M$-splines are piecewise polynomial functions of order $d$, which are defined on a set of $K+d$ knots, $t_{1},\ldots,t_{K+d}$. The basis function $M_{k}(y)$ is positive on $(t_{k},t_{k+d})$, and zero otherwise. Additionally $\int_{L}^{U}M_{k}(y)dy=1$, which gives $M_{k}(y)$ a natural interpretation as a density function supported on $[L,U]$. The basis functions are computed using the following system of recursive equations: for $d=1$,
@@
In the case of SPQR, we use an order of $d=3$ for the polynomial basis. We note that $B$-splines are also closely related to $M$-splines, as the $B$-spline basis function is defined by $B_{k}=(t_{k+d}-t_{k})M_{k}/d$ (Ramsay, 1988).
Ramsay (1988) also introduced integrated splines, or $I$-splines, over $[L,U]$. These are built using basis functions
@@
Appendix B Constructing valid SPQRx density functions
As discussed in Section 3.2, the distribution function defined in (4) is not guaranteed to be valid for arbitrary combinations of constituent distribution function $F(y|\mathcal{W})$, shape parameter $\xi$, and weight function $p(\cdot)$. When $F(y|\mathcal{W})$ is right-skewed and the constituent GP distribution function has relatively lighter-tails, and $F(y|\mathcal{W})>F_{\rm GP}(y|\mathcal{W})$ for some $y\in[a,b]$, we may find that the blended GP density function is negative for some $y\in[a,b]$. Whilst this is a pathological scenario that is unlikely to occur in practice (and did not affect the aforementioned simulation studies or application), we provide here a short discussion on this problem and practical advice for mitigating the risk of predicting invalid SPQRx distributions during model training.
Figure A1 gives an example of an invalid unconditional SPQRx distribution/density function. Here we have chosen $K=25$ monotonically-decreasing basis coefficients that ensure that $f_{\rm SPQR}$ is right-skewed on the interval $[0,1]$. We observe that the bulk of the $f_{\rm SPQR}$ density is at $y$ close to zero, with persistent non-zero mass $f_{\rm SPQR}(y)>0$ as $y$ goes towards its upper-endpoint at one. We take the constituent GP distribution to have shape parameter $\xi=-0.2$, choose the blending interval with levels $p_{a}=0.5$ and $p_{b}=0.99$, and set $c_{1}=c_{2}=5$ for the weighting function $p(y)$. As Figure A1 illustrates, we have $F_{\rm SPQR}(y|\mathcal{W})>F_{\rm GP}(y|\mathcal{W})$ for some $y\in[a,b],$ which results in the blended density $h(y|\mathcal{W},\xi)$ being negative for some $y$ in the blending interval. This is a pathological case that would not occur in practice: when $\xi$ and $f_{\rm SPQR}$ are jointly estimated, we would expect estimates of $\xi$ to increase to reflect the skewed behaviour of the $f_{\rm SPQR}$ density function.
Figure A2 illustrates three minor changes to the parameterisation of the distribution function in Figure A1 that will create a valid distribution. In the first example, we make the tails of $F_{\rm GP}$ heavier, by increasing $\xi$ to $0.1$. In the second example, we increase $c_{2}$, relative to $c_{1}$, which places more weight on the (valid) $F_{\rm SPQR}$ distribution function within the blending interval $[a,b]$. In the third example, we use a smaller blending interval, and lower $p_{b}$ from $0.99$ to $0.75$. These minor changes all lead to valid distribution functions.
Figure A1: Illustrative example of an invalid SPQRx distribution (left) and density (right) function. For $K=25$ specified basis coefficients, we construct the SPQR density and distribution functions (green curves). We then find $a=Q_{\rm SPQR}(p_{a}|\mathcal{W})$ and $b=Q_{\rm SPQR}(p_{b}|\mathcal{W})$, where here $p_{a}=0.5$ and $p_{b}=0.99$. The values of $a$ and $b$ are denoted by the orange horizontal lines, and the dashed horizontal line goes through zero. Then, we find the required GP distribution function (red) to satisfy continuity of the bGP($\mathcal{W},\xi)$ distribution function; its resulting distribution and density functions are provided in black. Here, we have $\xi=-0.2$ and $c_{1}=c_{2}=5$. Note that the blended GP density (black, right) is negative for some $y$.
Figure A2: Illustration of how the SPQRx distribution in Figure A1 can be adapted to create a valid distribution. Here, we change only one hyper-parameter, relative to Figure A1, per row. For the first row, we change $\xi$ from 0.2-0.2 to $0.1$. For the second row, we change $c_{1}$ from 5 to 100. For the third row, we change $p_{b}$ from 0.99 to 0.75. All SPQRx density functions (black) are positive for all $y$.
In a practical setting, one can mitigate the risk of predicting an invalid SPQRx distribution function during model training by: i) pre-training the SPQR density function and setting the initial $\xi$ estimate large (see Section 3.4); and ii) penalising invalid density functions. For the penalty, we add to the loss function
$\lambda\sum_{\mathbf{x}\in\{\mathbf{x}_{i}\}_{\rm train}}\int_{0}^{1}\text{ReLU}\left\{-h(y|\mathcal{W}(\mathbf{x}),\xi(\mathbf{x}))\right\}\mathrm{d}y,$
where $\{\mathbf{x}_{i}\}_{\rm train}$ denotes the training set of covariates, $\lambda>0$ is the penalty coefficient, and $\text{ReLU}(\cdot)=\max\{0,\cdot\}$ is the rectified linear unit activation function. In practice, we approximate the inner integral using a sequence of $y$ values on $[0,1]$. Throughout the main text and in Appendix C, we use both approaches during training, with $\lambda$ set to 100. In unreported experiments, we find that including the penalty improves the extrapolative performance of SQPRx, particularly when the shape parameter $\xi(\mathbf{x})$ is negative.
Appendix C Additional simulation studies
[…]
$\alpha_{0}(\mathbf{x})=3+\exp\{-1+\beta(\mathbf{x})\}>3,\quad\text{and}\quad\xi_{0}(\mathbf{x})=-1/[1+\exp\{-\beta(\mathbf{x})\}]\in(-1,0),$
with
$\beta_{0}(\mathbf{x})$ $=x_{2}x_{1}+x_{6}\left[1-\cos\left(\pi x_{3}x_{4}\right)\right]+2\sin(x_{5})/(|x_{7}-x_{8}|+2)$
$+0.2(x_{6}+x_{8}x_{9}/2)^{2}-\sqrt{x_{9}^{2}+x_{10}^{2}+2}.$
A similar function to $\beta_{0}$ was previously explored by Richards and Huser (2026b) and Richards and Huser (2026a). Note that, while $p=20$, only the first ten components of the covariate vector $\mathbf{X}$ act on the response distribution.
We conduct the two supplementary simulation studies as in Section 4. That is, over 100 repeated experiments, we evaluate the out-of-sample IWD and tIWD on a test set of 5000 observations, with models fitted using all considered candidate SPQR, SPQRx, and deep GP models, as described in Section 4. The results for settings 1 and 2 are presented in Tables A1 and A2, respectively. We note that, where the data are heavy-tailed (i.e., setting 1), the SPQRx model consistently outperforms SPQR, in terms of lower estimated IWD and tIWD. However, SPQRx provides lower tIWD than the deep GP model in the low sample size setting, where $n=1000$, suggesting that larger sample sizes may be required to see the benefits of extrapolation using SPQRx instead of standard peaks-over-threshold models. When the conditional upper-tails are bounded above (setting 2, Table A2), SPQRx is outperformed by SPQR. This is unsurprising as inference for SPQRx can be troublesome when the shape parameter is negative (see discussion in Section 3.4). Both SPQR and SPQRx are outperformed by the deep GP model, in terms of lower tIWD estimates. However, as discussed previously, the deep GP model is unable to provide a description for the full conditional density, and thus cannot provide IWD estimates.
Table A1: Results of the supplementary simulation study 1, with unit Lomax response distribution. The median of the IWD and tIWD estimates (alongside 50% confidence intervals; in brackets) are reported for the original SPQR, new SPQRx model (in bold), and the deep GP regression model (underlined). Note that the IWD is undefined for the deep GP model and the hyperparameter $K$ is not used in its construction. The hyper-parameters of the SPQRx, $(p_{a},p_{b},c_{1})$, are optimised for each row by minimising the tIWD, and are reported in the final column.
$n$ $K$ $n_{h}$ IWD tIWD Optimal $(p_{a},p_{b},c_{1})$
1000 15 16 0.052 (0.035, 0.075)/0.042 (0.029, 0.062) 3.76 (3.23, 5.27)/2.70 (1.46, 5.65) $(0.9,0.99,5)$
15 32 0.46 (0.034, 0.072)/0.040 (0.031, 0.065) 3.75 (3.20, 5.38)/3.17 (1.81, 5.94) $(0.75,0.99,25)$
25 16 0.048 (0.038, 0.063)/0.040 (0.030, 0.060) 3.76 (3.21, 5.41)/3.00 (2.14, 5.10) 2.47 (2.14, 3.14) $(0.9,0.99,25)$
25 32 0.045 (0.035, 0.057)/0.039 (0.028, 0.058) 3.76 (3.17, 5.37)/3.14 (2.00, 7.49) 2.46 (2.13, 3.09) $(0.75,0.99,25)$
10000 15 16 0.134 (0.078, 0.237)/0.132 (0.077, 0.235) 3.02 (2.26, 4.95)/2.22 (1.40, 3.45) $(0.75,0.999,5)$
15 32 0.134 (0.075, 0.238)/0.131 (0.076, 0.237) 2.90 (2.21, 4.94)/1.95 (1.22, 3.82) $(0.75,0.999,25)$
25 16 0.039 (0.025, 0.066)/0.037 (0.025, 0.069) 4.16 (3.08, 6.94)/1.40 ( 0.95, 2.71) 2.18 (2.02, 2.55) $(0.75,0.99,25)$
25 32 0.039 (0.027, 0.068)/0.037 (0.025, 0.070) 4.43 (2.89, 6.99)/1.33 (0.91, 2.51) 2.20 (1.94, 2.47) $(0.75,0.99,25)$
Table A2: Results of the supplementary simulation study 2, with a bounded GP response distribution. The median of the IWD and tIWD estimates (alongside 50% confidence intervals; in brackets) are reported for the original SPQR, new SPQRx model (in bold), and the deep GP regression model (underlined). Note that the IWD is undefined for the deep GP model and the hyperparameter $K$ is not used in its construction. The hyper-parameters of the SPQRx, $(p_{a},p_{b},c_{1})$, are optimised for each row by minimising the tIWD, and are reported in the final column.
$n$ $K$ $n_{h}$ IWD tIWD Optimal $(p_{a},p_{b},c_{1})$
1000 15 16 0.060 (0.056, 0.064)/0.059 (0.060, 0.063) 0.67 (0.59, 0.87)/0.65 (0.59, 0.80) $(0.75,0.99,5)$
15 32 0.059 (0.056, 0.064)/0.059 (0.056, 0.064) 0.67 (0.59, 0.90)/0.67 (0.59, 0.86) $(0.75,0.99,5)$
25 16 0.055 (0.053, 0.058)/0.055 (0.052, 0.058) 0.68 (0.60, 0.89)/0.66 (0.60, 0.88) 0.64 (0.56, 0.73) $(0.75,0.99,25)$
25 32 0.055 (0.052, 0.058)/0.054 (0.052, 0.056) 0.68 (0.60, 0.89)/0.66 (0.59, 0.89) 0.62 (0.55, 0.74) $(0.75,0.99,5)$
10000 15 16 0.061 (0.054, 0.064)/0.061 (0.054, 0.063) 0.66 (0.62, 0.75)/0.56 (0.50, 0.63) $(0.9,0.99,25)$
15 32 0.061 (0.054, 0.067)/0.060 (0.053, 0.063) 0.70 (0.65, 0.78)/0.59 (0.53, 0.65) $(0.75,0.99,25)$
25 16 0.047 (0.042, 0.051)/0.044 (0.042, 0.048) 0.73 (0.65, 0.82)/0.61 (0.57, 0.69) 0.41 (0.36, 0.44) $(0.75,0.99,25)$
25 32 0.046 (0.046, 0.051)/0.046 (0.043, 0.047) 0.74 (0.68, 0.81)/0.62 (0.57, 0.67) 0.33 (0.31, 0.35) $(0.75,0.99,25)$